19. Suppose that GDP (Y) is 5,000. Consumption is given by the consumption function C = 500 + 0.5(Y – T). Investment (I) is given by the investment function I = 2,000 – 100r, where r is the real interest rate in percent. Government spending (G) is 1,000 and net taxes (T) is also 1,000. When an increase in business optimism boosts the investment function to I = 3,000 – 100r: a. I rises by 1,000 and r rises by 10 percentage points. b. I rises by 1,000 and r is unchanged. c. I is unchanged and r rises by 10 percentage points. d. I is unchanged and r rises by 15 percentage points.
Answer #1

Answer : Investment rises by 1000 and r is unchanged.

Here intial investment = 2000-100r and new investment = 3000-100r. In the function, there is no change in slope, only the intercept changes. Intercept represents the autonomous investment. Autonomous investment is that investment which is independent of interest rate. Hence, autonomous investment increases by 1000.